Find an equation of the line having the given slope and containing the given point. Express your answerin the form x = a, y = b, or y = mx + b.
m = -7, (5,0)

Found 2 solutions by algebrahouse.com, mananth:
Answer by algebrahouse.com(1659)

Find an equation of the line having the given slope and containing the given point. Express your answerin the form x = a, y = b, or y = mx + b.
m = -7, (5,0)

Point-slope form can be used to get the equation of a line,
when given a point and a slope.
y - y1 = m(x - x1) is the point-slope form

(5,0), m = -7
y - 0 = -7(x - 5) {substituted into point-slope form}
y = -7x + 35 {used distributive property}

y = -7x + 35 is the equation of the line in "slope-intercept form"
7x + y = 35 is the equation of the line in "standard form"
www.algebrahouse.com

Answer by mananth(16949)

m= -7 ,point (5,0)
Find b by plugging the values of m & the point in
y=mx+b
0 =-35.00 +b
b= 35
m= 1.14
Plug value of the slope and b
The required equation is y=-7x+35